OpenStudy (anonymous):

I need help with finding the intercepts of this y=3(2- sqrt (x))/over x

OpenStudy (anonymous):

well y=0 u get x int/ and x=0 u get y int

OpenStudy (anonymous):

did it in my head so just check over

OpenStudy (anonymous):

but u get the idea

OpenStudy (anonymous):

actually nvm there's no y int

OpenStudy (anonymous):

i kinda need to c the part where your moving the square roots etcc

OpenStudy (anonymous):

beacuse its inverse function

OpenStudy (anonymous):

ok y=0

OpenStudy (anonymous):

0=3(2-sqrt(x))/x

OpenStudy (anonymous):

0=3(2-sqrt(x))

OpenStudy (anonymous):

0=(2-sqrt(x))

OpenStudy (anonymous):

-2=-sqrt(x)

OpenStudy (anonymous):

yes im following then you get 2-squrt(x)=0 which equals =4 then what hapens to he 3?

OpenStudy (anonymous):

3 gets divided out multiply both side by 1/3

OpenStudy (anonymous):

since anythin divided by 0 is 0 we can do tht

OpenStudy (anonymous):

ah so 1 x intercept, thanks!

OpenStudy (anonymous):

2=srt(x)

OpenStudy (anonymous):

4=x

OpenStudy (anonymous):

np